OpenStudy (anonymous):

The length of a rectangle is 3/8 its width. If its perimeter is 66, what is its length?

OpenStudy (anonymous):

Woops read the question wrong!

terenzreignz (terenzreignz):

Let's hope you're really not as username would suggest. First step in this sort of problem is representing values. We could let x be the length.

terenzreignz (terenzreignz):

If that's so... what's the width?

OpenStudy (anonymous):

This is a better drawing

OpenStudy (anonymous):

\[x+x+\frac{3x}{8}+\frac{3x}{8} = 66\]

OpenStudy (anonymous):

Thanks!! I've been stuck on this question for an hour already! :(